OpenStudy (anonymous):

x-c/2=d in terms of C show work please

OpenStudy (anonymous):

-c/2=d-x -> -c=2(d-x) -> c=-2(d-x)

OpenStudy (anonymous):

\[x-{c\over2}=d\]subtract x from both sides:\[-{c\over2}=d-x\]multiply both sides by -1 to remove the negative sign (and make life easier ;) ) :\[{c\over2}=-d+x~~~\implies~~~{c\over2}=x-d\]multiply both sides by 2 to remove the denominator:\[c={{x-d}\over2}\] hope this explanation helps! :)

OpenStudy (anonymous):

perfect thanks

OpenStudy (anonymous):

anytime! :D